# Spectral gap of a convex combination of a random permutation and a deterministic matrix

### v3

@@ 7.2 Encoding equivalent classes
and the multiplicity of an edge $e\in\mathrm{A}_{\gamma}$ is the number of $(i,t)$ verifying the previous equation. Then the set $\mathcal{Y}_{1}$ is in bijection with the corresponding incoming edges and these edges are inconsistent and of multiplicity one (see 6.5). Therefore we have $|\mathcal{Y}_{1}|\leq a_{1}$. If $y$ is seen by different edges, at least one of them is either discovered at a merging time or is an excess edge and therefore $|\mathcal{Y}_{3}|\leq g+m$. Finally the number of $y$ seen several times by the same edge and such that $|\mathcal{X}_{\operatorname{cc}}|\geq 2$ is bounded by the number of corresponding edges. We denote by $\mathrm{E}_{22}$ the number of edges $e:=(x,y)\in\mathrm{A}_{\gamma}$ linking $\operatorname{cc}^{\prime}$ to $\operatorname{cc}$ for some $\operatorname{cc}^{\prime},\operatorname{cc}\in\chi$ and such that $m(e)\geq 2$ and $|\mathcal{X}_{\operatorname{cc}}|\geq 2$. We proceed as in the proof of [BQZ18, Lemma 5] to bound $\mathrm{E}_{22}$. We call in-degree of a connected component $\operatorname{cc}\in\chi$ the number of edges $e\in\mathrm{A}_{\gamma}$ that links some $\operatorname{cc}^{\prime}\in\chi$ to $\operatorname{cc}$. Denoting $c_{k}$ the number of $\operatorname{cc}\in\chi$ with in-degree equal to $k$ and $c_{\geq 2}:=\sum_{k\geq 2}c_{k}$ we have:
$c_{0}+c_{1}+c_{\geq 2}=s-p\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ c_{1}+2c_{\geq 2}\leq\sum_{k}kc_{k}=a.$
@@
The bound $c_{0}\leq 2m$ follows from the observation that only a component $\operatorname{cc}\in\chi$ such that $\operatorname{cc}=\operatorname{cc}(x_{j,1})$ for some $1\leq j\leq 2m$ can be of in-degree 00. Besides, denoting $\mathrm{E}_{12}$ the number of edges of $\mathrm{A}_{\gamma}$ that are of multiplicity at least $2$ and such that the outgoing connected component $\operatorname{cc}\in\chi$ verifies $|\mathcal{X}_{\operatorname{cc}}|=1$, we observe that $c_{1}\leq a_{1}+\mathrm{E}_{12}$. Indeed we can associate each connected component $\operatorname{cc}$ of in-coming degree equal to one to its only in-coming edge. In particular we have
$|\mathcal{Y}_{3}|\leq\mathrm{E}_{22}=a-a_{1}-\mathrm{E}_{12}\leq a-c_{1}\leq 2g+2m-2\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and therefore}\leavevmode\nobreak\ \leavevmode\nobreak\ F\leq 3(m+g)-2+a_{1}+p.$
The last lemma of this section will estimate the contribution $\mu(\gamma)$ defined in (28) using 6.5.
@@
There is a constant $c>0$ such that, if $\gamma\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)$, $g=a-s+p+1$ and $a_{1}$ is the number of arcs in $\mathrm{A}_{\gamma}$ visited exactly once by $\gamma$ then
@@
Let $\mathrm{A}_{\ast}$ be the set of inconsistent arcs (see definitions above 6.5) and we denote $b$ its cardinal. We consider $\mathrm{A}_{1}^{\prime}\subset\mathrm{A}_{1}$ the subset of consistent arcs of multiplicity one. We have
@@
$a_{1}+a_{\geq 2}=a\leavevmode\nobreak\ \leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \leavevmode\nobreak\ a_{1}+2a_{\geq 2}\leq\sum_{k}ka_{k}=T\leq 2m(\ell+1).$
@@
It gives the second claim of the lemma. We now count the inconsistent edges using the vocabulary of the the proof of 7.2. Exploring $\gamma$ chronologically, seeing a new inconsistent edge means that we discover an edge with either one of its ending vertices that had been seen before. That can happen either at the beginning of a sequence of times that goes out of the existing forest so far, at a merging time or at a first visit of an excess edge. Every time $(i,t)$ where we start a sequence of times going out of the existing forest and such that we discover a new inconsistent edge is preceded by a first visit of an excess edge. Every such step can create $2$ inconsistent edge: the one we see at this time and the one seen before with at least one of its ending point in common. Therefore we have
@@
It remains to apply 6.5. ∎
We have all the ingredients to now prove 7.1.
Proof of 7.1.
We prove the case #=#\#$\emptyset$# = ∅ and all the computations are exactly the same for #$\#=r$. We set
@@
$\mathbb{P}\left(\|\underline{A}^{(\ell)}\|\geq e^{C\ln(N)^{1-\alpha}}\rho_{n}^{\ell}\right)\leq\mathbb{P}\left(\|\underline{A}^{(\ell)}\|^{2m}\geq N^{C}\rho_{n}^{2m\ell}\right).$
@@
$S:=\sum_{s,a,p,\vec{T}}\left|\mathcal{W}_{\ell,\vec{T}}(s,a,p)\right|\underset{\gamma\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)}{\max}\left\{|\mu(\gamma)|\sum_{\gamma^{\prime}}q(\gamma^{\prime})\right\}\leq N^{\mathfrak{c}}\rho_{n}^{2m\ell}.$ (34)
Let $\gamma\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)$ and $a_{1}$ be the number of $(x,y)\in\mathrm{A}_{\gamma}$ of multiplicity one. Applying 7.5 and 7.9 we have
$|\mu(\gamma)|\sum_{\gamma^{\prime}\sim\gamma}q(\gamma^{\prime})\leq N^{s-p}(cm\ell)^{a_{1}+3(m+g)+p}c(n)\rho_{n}^{2m(\ell+1)}c^{m+g}N^{-a}\left(\frac{6\ell m}{\sqrt{N}}\right)^{(a_{1}-4g-2m+2p)_{+}}.$
In regard of the conditions on $Q$ in 2.2 and the definition of $\rho_{n}$, $\ell$ and $m$, we find $c>1$ such that $\rho_{n}\leq c\ell m$. Since $cm\ell\geq 1$ we have $(cm\ell)^{a_{1}+p}\leq(cm\ell)^{(a_{1}-4g-2m+2p)_{+}}(2m\ell)^{4g+2m}$ and using that $a_{1}\geq 2(a-m(\ell+1))$ we deduce:
$|\mu(\gamma)|\sum_{\gamma^{\prime}\sim\gamma}q(\gamma^{\prime})\leq c(n)(cm\ell)^{8g+8m}N^{-g+1}\rho_{n}^{2m\ell}\left(\frac{(cm\ell)^{4}}{N}\right)^{(a-(\ell+2)m-2g+p)_{+}},$
@@
Considering any $\ell=\mathcal{O}(\ln(N))$, we have $\epsilon=o(1)$. We apply 7.2 to $S$ defined by (34) and sum over all $\vec{T}$ possible. We have at most $\ell^{2m}$ possibilities for $\vec{T}$ and considering the fact that for $c>1$ large enough we have $a,s$ and $\ell$ bounded by $cm\ell$ we obtain
$S^{\prime}$ $:=\frac{S}{\rho_{n}^{2m\ell}}\leq c(n)N\sum_{\begin{subarray}{c}s,a,p\\ g(s,a,p)\geq 0,\leavevmode\nobreak\ p\leq 2(m+g)\end{subarray}}\ell^{2m}(2m\ell)^{12m(g+5)}(cm\ell)^{8g+8m}N^{-g}\epsilon^{(a-(\ell+2)m-2g+p)_{+}}$
$\leq cN(cm\ell)^{80m}\sum_{\begin{subarray}{c}s,g,p\\ g(s,a,p)\geq 0,\leavevmode\nobreak\ p\leq 2(m+g)\end{subarray}}(cm\ell)^{20mg}N^{-g}\epsilon^{(s-\ell^{\prime}m-g)_{+}},$
where we performed the change of variable $g=a-s+p+1$ and where we consider $\ell^{\prime}:=\ell+2+1/m$. Now summing over $p\leq 2(m+g)$ we have in the sum an additional term $2(m+g)$ which we bound with $cm\ell$. Considering $\mathfrak{c}>0$ such that $c(n)(cm\ell)^{81m}\leq N^{\mathfrak{c}}$:
@@
7.4 Operator norm of $R^{(\ell)}_{k,\#}$
We will now adapt what have been doing to bound the operator norm $A_{\#}^{(\ell)}$ to upper bound the operator norm of $R^{(\ell)}_{k,\#}$ defined by (24) for #=#\#$\emptyset$# = ∅ and (25) for #$\#=r$. We set
$\Delta:=1+\kappa+\delta\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \Delta_{r}:=1+r\kappa+r\delta$
@@ 7.4 Operator norm of $R^{(\ell)}_{k,\#}$
Let $(Q,M),\leavevmode\nobreak\ n,\leavevmode\nobreak\ \rho_{n}$ and $A$ be as defined in 2.2. For any $c>0$, $10<\alpha<$ 1 and $1\leq k\leq\ell\leq\frac{\ln(N)}{7\ln(d)}$, there exists $C=C(c,\alpha,n,\delta)>0$ such that with probability at least $1-N^{-c}$,
$\|R_{k,\#}^{(\ell)}\|\leq e^{C\ln(N)^{1-\alpha}}\Delta_{\#}^{\ell}.$
In all this section the notations are the one introduced in Section 7.1. We also give the proof of the previous proposition only in the case #=#\#$\emptyset$# = ∅. We again use the inequality
$\|R^{(\ell)}_{k}\|^{2m}\leq\operatorname{Tr}\left\{\left(R_{k}^{(\ell)}{R_{k}^{(\ell)}}^{\top}\right)^{m}\right\}=:\operatorname{Tr}[k,\ell,m].$ (35)
@@
As in Section 7.1 we define the graph $K_{\gamma}$ for a fixed $\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}$. We consider $X_{\gamma}:=\{x_{i,t}:i\in[2m],t\in[T_{i}+1]\}$ and $Y_{\gamma}:=\{y_{i,t}:i\in[2m],t\in[T_{i}]\}$. As before we set $K_{\gamma}=(X_{\gamma},E(K_{\gamma}))$ and $\{x,x^{\prime}\}$ is a vertex if there exists $(i,t),(j,s)\in\boldsymbol{T}$ such that
$x_{i,t+1}=x,\leavevmode\nobreak\ x_{j,s+1}=x^{\prime}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \leavevmode\nobreak\ y_{i,t}=y_{j,s}.$
For $x\in X_{\gamma}$ we consider again $\operatorname{cc}(x)$ its connected component in $K_{\gamma}$ and $\chi_{\gamma}=X_{\gamma}/\sim$ the set of connected component of $K_{\gamma}$. The arcs of $\gamma$ are the pairs $(x_{i,t},y_{i,t})$ for $t\neq k$ and we denote by $\mathrm{A}^{\prime}_{\gamma}$ the set of arcs of $\gamma$. For $s,a,p$ integers, we denote by $\mathrm{W}^{\prime}_{\ell,\vec{T}}(s,a,p)$ the set of $\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}$ verifying $|X_{\gamma}|=s$, $|X_{\gamma}/\sim|=s-p$ and $|\mathrm{A}_{\gamma}^{\prime}|=a$. Adapting the computations for the case #$\#=r$ and taking the expectation of (36) we have
$\mathbb{E}\|R_{k,\#}^{(\ell)}\|^{2m}\leq\mathbb{E}(\operatorname{Tr}[k,\ell,m])=\sum_{\vec{T},s,a,p}\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}(s,a,p)}\mu^{\prime}(\gamma)q_{\#}(\gamma)$
@@
$q(\gamma):=\prod_{i=1}^{2m}\prod_{t=1}^{T_{i}}(Q^{\ell_{i,t}})_{y_{i,t}x_{i,t+1}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \leavevmode\nobreak\ q_{r}(\gamma):=(1-r)^{T-1}\prod_{i=1}^{2m}\prod_{t=1}^{T_{i}}(\{rQ\}^{\ell_{i,t}})_{y_{i,t}x_{i,t+1}}$ (38)
@@
$\mathbb{E}\|R_{k,\#}^{(\ell)}\|^{2m}\leq\sum_{\vec{T},s,a,p}|\mathcal{W}^{\prime}_{\ell,\vec{T}}(s,a,p)|\underset{\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}(s,a,p)}{\max}\left(|\mu^{\prime}(\gamma)|\sum_{\gamma^{\prime}\sim\gamma}q_{\#}(\gamma^{\prime})\right).$ (39)
@@
We repeat the construction given in Section 7.2. Let $\ell\geq 1$, $\vec{T}\in[\ell]^{2m}$ and $\gamma\in\mathrm{W}^{\ell,\vec{T}}(s,a,p)$ be fixed. We set $\gamma_{i,t}:=(x_{i,t},y_{i,t},x_{i,t+1})$ for $i\in[2m]$ and $t\in[T_{i}]$. We explore $(\gamma_{i,t})_{(i,t)\in\boldsymbol{T}}$ chronologically where $\boldsymbol{T}:=\{(i,t):i\in[2m],t\in[T_{i}]\backslash\{k\}\}$ is ordered as in Section 7.2. Recall in particular that for $i\leq i^{\prime}$, for all $t\in[T_{i}]\backslash\{k\}$ and $t^{\prime}\in[T_{i^{\prime}}]\backslash\{k\}$ we have $(i,t)\preceq(i^{\prime},t^{\prime})$, and for all $t\leq t^{\prime}\in[T_{i}]\backslash\{k\}$ we have $(i,t)\preceq(i,t^{\prime})$.
@@
For each even era $i\in[2m]$, we define the first merging time $(i,t^{\prime}_{i})$ as the smallest time $(i,t)$ with $11\leq t\leq k-$ 1, if it exists, such that $\Gamma_{i,t}$ and $\Gamma_{(i,1)^{-}}$ have the same number of connected component. If this time does not exists exist we take as convention $t^{\prime}_{i}=k$. Similarly, for each $i$, we define the second merging time $(i,t_{i}^{\operatorname{\prime\prime}})$ as the smallest time $(i,t)$ with $k\leq t\leq T_{i}$ such that $\Gamma_{i,t}$ has the same number of connected components than $\Gamma_{(i,k)^{-}}$. If this time does not exists we set $t^{\operatorname{\prime\prime}}_{i}=T_{i}+1$. In regard of (26), we have $t^{\operatorname{\prime\prime}}_{i}\leq T_{i}$ for $i$ even. Also notice that if $t_{i}^{\epsilon}\geq 2$ then the merging time is a first time.

### v4

@@ 8.2 Encoding equivalent classes
and the multiplicity of an edge $e\in\mathrm{A}_{\gamma}$ is the number of $(i,t)$ verifying the previous equation. Then the set $\mathcal{Y}_{1}$ is in bijection with the corresponding incoming edges and these edges are inconsistent and of multiplicity one (see Section˜7.3). Therefore we have $|\mathcal{Y}_{1}|\leq a_{1}$. If $y$ is seen by different edges, at least one of them is either discovered at a merging time or is an excess edge and therefore $|\mathcal{Y}_{3}|\leq g+m$. Finally the number of $y$ seen several times by the same edge and such that $|\mathcal{X}_{\operatorname{cc}}|\geq 2$ is bounded by the number of corresponding edges. We denote by $\mathrm{E}_{22}$ the number of edges $e:=(x,y)\in\mathrm{A}_{\gamma}$ linking $\operatorname{cc}^{\prime}$ to $\operatorname{cc}$ for some $\operatorname{cc}^{\prime},\operatorname{cc}\in\chi$ and such that $m(e)\geq 2$ and $|\mathcal{X}_{\operatorname{cc}}|\geq 2$. We proceed as in the proof of [BQZ18, Lemma 5] to bound $\mathrm{E}_{22}$. We call in-degree of a connected component $\operatorname{cc}\in\chi$ the number of edges $e\in\mathrm{A}_{\gamma}$ that links some $\operatorname{cc}^{\prime}\in\chi$ to $\operatorname{cc}$. Denoting $c_{k}$ the number of $\operatorname{cc}\in\chi$ with in-degree equal to $k$ and $c_{\geq 2}:=\sum_{k\geq 2}c_{k}$ we have:
$c_{0}+c_{1}+c_{\geq 2}=s-p\penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ c_{1}+2c_{\geq 2}\leq\sum_{k}kc_{k}=a.$
@@
The bound $c_{0}\leq 2m$ follows from the observation that only a component $\operatorname{cc}\in\chi$ such that $\operatorname{cc}=\operatorname{cc}(x_{j,1})$ for some $1\leq j\leq 2m$ can be of in-degree 0. Besides, denoting $\mathrm{E}_{12}$ the number of edges of $\mathrm{A}_{\gamma}$ that are of multiplicity at least $2$ and such that the outgoing connected component $\operatorname{cc}\in\chi$ verifies $|\mathcal{X}_{\operatorname{cc}}|=1$, we observe that $c_{1}\leq a_{1}+\mathrm{E}_{12}$. Indeed we can associate each connected component $\operatorname{cc}$ of in-coming degree equal to one to its only in-coming edge. In particular we have
$|\mathcal{Y}_{3}|\leq\mathrm{E}_{22}=a-a_{1}-\mathrm{E}_{12}\leq a-c_{1}\leq 2g+2m-2\penalty 10000\ \penalty 10000\ \text{and therefore}\penalty 10000\ \penalty 10000\ F\leq 3(m+g)-2+a_{1}+p.$
Remark 8.2.
One strategy to obtain Section˜3 this conjecture would be to get an analog of [BQZ18, Lemma 4] in the following sense. We would need to improve the combinatorial arguments at stake at stake in ˜7.2 and ˜8.6 in order to obtain that when summing over all $\gamma^{\prime}\sim\gamma$ for some $\vec{\ell}$ fixed tangle-free in (36), we obtain
$\sum_{\begin{subarray}{c}\gamma^{\prime}\sim\gamma\\ \vec{\ell}(\gamma^{\prime})=\vec{\ell}\end{subarray}}q_{\mathrm{s}}(\gamma^{\prime})\leq N^{s-p}\delta^{L_{\max}}\kappa^{L_{\mathcal{R}}}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(Q^{\ell(e(\operatorname{cc}))}{Q^{\ast}}^{\ell(e^{\prime}(\operatorname{cc}))}\right)$
for some $L_{\max}\leq CmF$ with $C>0$ a universal constant.
The last lemma of this section will estimate the contribution $\mu(\gamma)$ defined in (31) using Section˜7.3.
@@
There is a constant $c>0$ such that, if $\gamma\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)$, g=as+p+1g=a-s+p+1 and $a_{1}$ is the number of arcs in $\mathrm{A}_{\gamma}$ visited exactly once by $\gamma$ then
@@
Let $\mathrm{A}_{\ast}$ be the set of inconsistent arcs (see definitions above Section˜7.3) and we denote $b$ its cardinal. We consider $\mathrm{A}_{1}^{\prime}\subset\mathrm{A}_{1}$ the subset of consistent arcs of multiplicity one. We have
@@
$a_{1}+a_{\geq 2}=a\penalty 10000\ \penalty 10000\ \text{ and }\penalty 10000\ \penalty 10000\ a_{1}+2a_{\geq 2}\leq\sum_{k}ka_{k}=T\leq 2m(\ell+1).$
@@
It gives the second claim of the lemma. We now count the inconsistent edges using the vocabulary of the the proof of 8.1. Exploring $\gamma$ chronologically, seeing a new inconsistent edge means that we discover an edge with either one of its ending vertices that had been seen before. That can happen either at the beginning of a sequence of times that goes out of the existing forest so far, at a merging time or at a first visit of an excess edge. Every time $(i,t)$ where we start a sequence of times going out of the existing forest and such that we discover a new inconsistent edge is preceded by a first visit of an excess edge. Every such step can create $2$ inconsistent edge: the one we see at this time and the one seen before with at least one of its ending point in common. Therefore we have
@@
It remains to apply Section˜7.3. ∎
We have all the ingredients to now prove Section˜8.1.
Proof of Section˜8.1.
We prove the case $\mathrm{x}=\mathrm{s}$ and all the computations are exactly the same for $\mathrm{x}=\mathrm{s}$. We set
@@
$\mathbb{P}\left(\|\underline{A}_{\mathrm{s}}^{(\ell)}\|\geq e^{C\ln(N)^{1-\alpha}}\rho_{n,\mathrm{s}}^{\ell}\right)\leq\mathbb{P}\left(\|\underline{A}_{\mathrm{s}}^{(\ell)}\|^{2m}\geq N^{C}\rho_{n,\mathrm{s}}^{2m\ell}\right).$
@@
$S:=\sum_{s,a,p,\vec{T}}\left|\mathcal{W}_{\ell,\vec{T}}(s,a,p)\right|\underset{\gamma\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)}{\max}\left\{|\mu(\gamma)|\sum_{\gamma^{\prime}}q(\gamma^{\prime})\right\}\leq N^{\mathfrak{c}}\rho_{n,\mathrm{s}}^{2m\ell}.$ (42)
Let $\gamma\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)$ and $a_{1}$ be the number of $(x,y)\in\mathrm{A}_{\gamma}$ of multiplicity one. Applying ˜8.4 and ˜8.7 we have
$|\mu(\gamma)|\sum_{\gamma^{\prime}\sim\gamma}q_{\mathrm{s}}(\gamma^{\prime})\leq N^{s-p}(cm\ell)^{a_{1}+3(m+g)+p}c(n)\rho_{n,\mathrm{s}}^{2m(\ell+1)}c^{m+g}N^{-a}\left(\frac{6\ell m}{\sqrt{N}}\right)^{(a_{1}-4g-2m+2p)_{+}}.$
In regard of the conditions on $Q$ in ˜2.1 and the definition of $\rho_{n}$, $\ell$ and $m$, we find $c>1$ such that $\rho_{n}\leq c\ell m$. Since $cm\ell\geq 1$ we have $(cm\ell)^{a_{1}+p}\leq(cm\ell)^{(a_{1}-4g-2m+2p)_{+}}(2m\ell)^{4g+2m}$ and using that $a_{1}\geq 2(a-m(\ell+1))$ we deduce:
$|\mu(\gamma)|\sum_{\gamma^{\prime}\sim\gamma}q_{\mathrm{s}}(\gamma^{\prime})\leq c(n)(cm\ell)^{8g+8m}N^{-g+1}\rho_{n,\mathrm{s}}^{2m\ell}\left(\frac{(cm\ell)^{4}}{N}\right)^{(a-(\ell+2)m-2g+p)_{+}},$
@@
Considering any $\ell=\mathcal{O}(\ln(N))$, we have $\epsilon=o(1)$. We apply ˜8.1 to $S$ defined by (42) and sum over all $\vec{T}$ possible. We have at most $\ell^{2m}$ possibilities for $\vec{T}$ and considering the fact that for $c>1$ large enough we have $a,s$ and $\ell$ bounded by $cm\ell$ we obtain
$S^{\prime}$ $:=\frac{S}{\rho_{n,\mathrm{s}}^{2m\ell}}\leq c(n)N\sum_{\begin{subarray}{c}s,a,p\\ g(s,a,p)\geq 0,\penalty 10000\ p\leq 2(m+g)\end{subarray}}\ell^{2m}(2m\ell)^{12m(g+5)}(cm\ell)^{8g+8m}N^{-g}\epsilon^{(a-(\ell+2)m-2g+p)_{+}}$
$\leq cN(cm\ell)^{80m}\sum_{\begin{subarray}{c}s,g,p\\ g(s,a,p)\geq 0,\penalty 10000\ p\leq 2(m+g)\end{subarray}}(cm\ell)^{20mg}N^{-g}\epsilon^{(s-\ell^{\prime}m-g)_{+}},$
where we performed the change of variable g=as+p+1g=a-s+p+1 and where we consider $\ell^{\prime}:=\ell+2+1/m$. Now summing over $p\leq 2(m+g)$ we have in the sum an additional term $2(m+g)$ which we bound with $cm\ell$. Considering $\mathfrak{c}>0$ such that $c(n)(cm\ell)^{81m}\leq N^{\mathfrak{c}}$:
@@
8.4 Operator norm of $R^{(\ell)}_{k,\mathrm{x}}$
We will now adapt what have been doing to bound the operator norm $A_{\mathrm{x}}^{(\ell)}$ to upper bound the operator norm of $R^{(\ell)}_{k,\mathrm{x}}$ defined by (27) for $\mathrm{x}=\mathrm{s}$ and (28) for $\mathrm{x}=\mathrm{c}$. We set
$\Delta_{\mathrm{s}}:=1+\kappa+\delta\penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ \Delta_{\mathrm{c}}:=1+r\kappa+r\delta$
@@ 8.4 Operator norm of $R^{(\ell)}_{k,\mathrm{x}}$
Let $(Q,M),\penalty 10000\ n,\penalty 10000\ \rho_{n}$ and $A$ be as defined in ˜2.1. For any $c>0$, $0<\alpha<1$ and $1\leq k\leq\ell\leq\frac{\ln(N)}{7\ln(d)}$, there exists $C=C(c,\alpha,n,\delta)>0$ such that with probability at least $1-N^{-c}$,
$\|R_{k,\mathrm{x}}^{(\ell)}\|\leq e^{C\ln(N)^{1-\alpha}}\Delta_{\mathrm{x}}^{\ell}.$
In all this section the notations are the one introduced in Section˜8.1. We also give the proof of the previous proposition only in the case $\mathrm{x}=\mathrm{s}$. We again use the inequality
$\|R^{(\ell)}_{k,\mathrm{s}}\|^{2m}\leq\operatorname{Tr}\left\{\left(R_{k,\mathrm{s}}^{(\ell)}{R_{k,\mathrm{s}}^{(\ell)}}^{\top}\right)^{m}\right\}=:\operatorname{Tr}[k,\ell,m].$ (43)
@@
As in Section˜8.1 we define the graph $K_{\gamma}$ for a fixed $\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}$. We consider $X_{\gamma}:=\{x_{i,t}:i\in[2m],t\in[T_{i}+1]\}$ and $Y_{\gamma}:=\{y_{i,t}:i\in[2m],t\in[T_{i}]\}$. As before we set $K_{\gamma}=(X_{\gamma},E(K_{\gamma}))$ and $\{x,x^{\prime}\}$ is a vertex if there exists $(i,t),(j,s)\in\boldsymbol{T}$ such that
$x_{i,t+1}=x,\penalty 10000\ x_{j,s+1}=x^{\prime}\penalty 10000\ \penalty 10000\ \text{ and }\penalty 10000\ \penalty 10000\ y_{i,t}=y_{j,s}.$
For $x\in X_{\gamma}$ we consider again $\operatorname{cc}(x)$ its connected component in $K_{\gamma}$ and $\chi_{\gamma}=X_{\gamma}/\sim$ the set of connected component of $K_{\gamma}$. The arcs of $\gamma$ are the pairs $(x_{i,t},y_{i,t})$ for $t\neq k$ and we denote by $\mathrm{A}^{\prime}_{\gamma}$ the set of arcs of $\gamma$. For $s,a,p$ integers, we denote by $\mathrm{W}^{\prime}_{\ell,\vec{T}}(s,a,p)$ the set of $\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}$ verifying $|X_{\gamma}|=s$, $|X_{\gamma}/\sim|=s-p$ and $|\mathrm{A}_{\gamma}^{\prime}|=a$. Adapting the computations for the case $\mathrm{x}=\mathrm{c}$ and taking the expectation of (44) we have
$\mathbb{E}\|R_{k,\mathrm{x}}^{(\ell)}\|^{2m}\leq\mathbb{E}(\operatorname{Tr}[k,\ell,m])=\sum_{\vec{T},s,a,p}\penalty 10000\ \penalty 10000\ \sum_{\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}(s,a,p)}\mu^{\prime}(\gamma)q_{\mathrm{x}}(\gamma)$
@@
$q_{\mathrm{s}}(\gamma):=\prod_{i=1}^{2m}\prod_{t=1}^{T_{i}}(Q^{\ell_{i,t}})_{y_{i,t}x_{i,t+1}}\penalty 10000\ \penalty 10000\ \text{ and }\penalty 10000\ \penalty 10000\ q_{\mathrm{c}}(\gamma):=(1-r)^{T-1}\prod_{i=1}^{2m}\prod_{t=1}^{T_{i}}(\{rQ\}^{\ell_{i,t}})_{y_{i,t}x_{i,t+1}}$ (46)
@@
$\mathbb{E}\|R_{k,\mathrm{x}}^{(\ell)}\|^{2m}\leq\sum_{\vec{T},s,a,p}|\mathcal{W}^{\prime}_{\ell,\vec{T}}(s,a,p)|\underset{\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}(s,a,p)}{\max}\left(|\mu^{\prime}(\gamma)|\sum_{\gamma^{\prime}\sim\gamma}q_{\mathrm{x}}(\gamma^{\prime})\right).$ (47)
@@
[…]
@@
For each even era $i\in[2m]$, we define the first merging time $(i,t^{\prime}_{i})$ as the smallest time $(i,t)$ with $1\leq t\leq k-1$, if it exists, such that $\Gamma_{i,t}$ and $\Gamma_{(i,1)^{-}}$ have the same number of connected component. If this time does not exists exist we take as convention $t^{\prime}_{i}=k$. Similarly, for each $i$, we define the second merging time $(i,t_{i}^{\operatorname{\prime\prime}})$ as the smallest time $(i,t)$ with $k\leq t\leq T_{i}$ such that $\Gamma_{i,t}$ has the same number of connected components than $\Gamma_{(i,k)^{-}}$. If this time does not exists we set $t^{\operatorname{\prime\prime}}_{i}=T_{i}+1$. In regard of (29), we have $t^{\operatorname{\prime\prime}}_{i}\leq T_{i}$ for $i$ even. Also notice that if $t_{i}^{\epsilon}\geq 2$ then the merging time is a first time.